OpenStudy (shuckle):

Given 6(y+3) = 2(x+9) find the ration of x to y?

OpenStudy (mathstudent55):

First, distribute the 6 on the left side and the 2 on the right side. Can you do that and show it?

OpenStudy (shuckle):

6y +18 = 2x + 18

OpenStudy (mathstudent55):

Great. Now subtract 18 from both sides.

OpenStudy (shuckle):

6y = 2x?

OpenStudy (mathstudent55):

Correct. You have 6y = 2x which is the same as 2x = 6y

OpenStudy (mathstudent55):

You want x/y = something

OpenStudy (mathstudent55):

Now divide both sides by 2 starting with 2x = 6y

OpenStudy (shuckle):

2x/6y + 1:3?

OpenStudy (shuckle):

= I mean

OpenStudy (mathstudent55):

Do what I wrote above. Start with 2x = 6y and divide both sides by 2.

OpenStudy (shuckle):

1x= 3y

OpenStudy (mathstudent55):

Great. Almost there. Since we want x/y on the left side, now just divide both sides by y.

OpenStudy (shuckle):

1x/y = 3?

OpenStudy (mathstudent55):

Yes. 1x/y is simply x/y, so you can just write x/y on the left side. The right side is simply 3. Since you are trying to write what ratio the left side is equal to, you can write the right side as 3/1 since 3 = 3/1 Then you have: x/y = 3/1 In other words, the ratio of x to y is 3 to 1.

OpenStudy (shuckle):

Thank you. :)

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (triciaal):